OpenStudy (anonymous):

A sound wave is modeled with the equation y=1/4cos2pi/3theta a. Find the period. Explain your method. b. Find the amplitude. Explain your method.

OpenStudy (campbell_st):

given a general equation cos(bx) the period is \[\frac{2\pi}{b}\] in your equation the value of b is \[b = \frac{2\pi}{3}\] substitute it to find the period of the curve

OpenStudy (anonymous):

wait so would it be 2.09?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well its \[period = (2\pi) \div \frac{2\pi}{3}\]

OpenStudy (anonymous):

I put it in the calc and it says 3.

OpenStudy (campbell_st):

well thats the period.... the curve completes 1 full cycle over a distance of 3 units. now the amplitude the general form is y =acos(bx) a is the amplitude... which is furthest distance the curve moves from its centre. so looking at your curve y = 1/4cos(2pi/3 x) what do you think the amplitude is?

OpenStudy (anonymous):

uhh.. I'm so bad at this I don't know.. -1/8?? Idk @campbell_st

OpenStudy (campbell_st):

a = 1/4 so thats the amplitude..... amplitude is always the number infront of the trig ratio... if no number is written the amplitude is 1.

OpenStudy (anonymous):

Ohhh yeah, that's right. Thanks so much for your help!